How many cubic feet of cement are in a rectangular cement slab that is 3 inlches thick 8.5 inches long and 12 feet wide?
Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

How many cubic feet of cement are in a rectangular cement slab that is 3 inlches thick 8.5 inches long and 12 feet wide?
------------
12 feet = 144 inches
Vol = 3*8.5*144 cubic inches = 3672
1 cubic foot = 1728 cubic inches (12^3)
Vol = 3672/1728 cubic feet
Vol = 2.125 cubic ft

Answer by edjones(8007)

3/12 * 8.5/12 * 12 = 306/144 = 2.125 cu.ft.
.
Ed